OpenStudy (anonymous):

Solve using substitution 2x=y+9 5y=10x-45

OpenStudy (lgbasallote):

if 2x = y + 9 then let y = 2x - 9 make sense?

OpenStudy (anonymous):

no really :/

OpenStudy (anonymous):

find one variable, either x or y, first

OpenStudy (anonymous):

i think ill choose x

OpenStudy (anonymous):

so if you get x alone for one equation, you then replace what you get into the second equation and solve for y

OpenStudy (anonymous):

so i mutiply 10x on all 3 of the first equation?

OpenStudy (anonymous):

o.O? to get x alone? or?

OpenStudy (anonymous):

to subtract it with the other equation?

OpenStudy (anonymous):

are you trying to solve it by elimination?

OpenStudy (anonymous):

no by substitution

OpenStudy (anonymous):

if you get Y alone in the first equation it would be what the first person said y=2x-9 then you substitute the 2x-9 into the second equation getting 5(2x-9)=10x+45

OpenStudy (anonymous):

10x-45=10x-45*

OpenStudy (anonymous):

oh i kinda get it. then you multipy 5 on the parantesis

OpenStudy (anonymous):

than you add like terms?

OpenStudy (anonymous):

both sides equal each other if you combine like terms you get 0=0

OpenStudy (anonymous):

:o o? wow. now i have to solve the other 1

OpenStudy (anonymous):

good luck! hope it helped.

OpenStudy (anonymous):

its sure did friend=)